# Your AI, My Shell”: Demystifying Prompt Injection Attacks on Agentic AI Coding Editors

### v1

@@ Abstract.
Agentic AI coding editors driven by large language models have recently become more popular due to their ability to improve developer productivity during software development. Modern editors such as Cursor are designed not just for code completion, but also with more system privileges for complex coding tasks (e.g., run commands in the terminal, access development environments, and interact with external systems). While this brings us closer to the “fully automated programming” dream, it also raises new security concerns. In this study, we present the first empirical analysis of prompt injection attacks targeting these high-privilege agentic AI coding editors. We show how attackers can remotely exploit these systems by poisoning external development resources with malicious instructions, effectively hijacking AI agents to run malicious commands, turning “your AI” into “attacker’s shell”. To perform this analysis, we implement AIShellJack, an automated testing framework for assessing prompt injection vulnerabilities in agentic AI coding editors. AIShellJack  contains 314 unique attack payloads that cover 70 techniques from the MITRE ATT&CK framework. Using AIShellJack, we conduct a large-scale evaluation on GitHub Copilot and Cursor, and our evaluation results show that attack success rates can reach as high as 84% for executing malicious commands. Moreover, these attacks are proven effective across a wide range of objectives, ranging from initial access and system discovery to credential theft and data exfiltration.
Prompt Injection, Large Language Models, AI Security, Agentic AI
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: PACMSEjournalvolume: 1journalnumber: 1article: 1publicationmonth: 10ccs: Security and privacy Software security
1. Introduction
@@
Powered by breakthroughs in large language models (LLMs), AI pair-programming tools promise to change the face of software development. Early tools like GitHub Copilot (GitHub, 2025a) and Amazon Q Developer (Amazon, 2025) improve developer productivity by providing real-time code suggestions, and autocompletions. There is no doubt that these tools have greatly improved developer productivity and obtained significant adoption (Ziegler et al., 2024; Lyu et al., 2025). A survey by Stack Overflow shows that software developers are increasingly using or plan to use AI pair-programming tools, a considerable increase from 76% in 2024 to 84% in 2025 (Stack Overflow, 2025), with 51% of professionals using them daily. Thanks to recent advances in LLMs, AI pair programmers is now deeply integrated into our development environments and can autonomously plan and execute complex coding tasks, which we call agentic AI coding editors. GitHub Copilot coding agent is now deeply integrated with the latest VSCode releases, which can carry out coding tasks autonomously in the background (e.g., manage pull requests) (GitHub, 2025b). Cursor (Cursor, 2025b), another popular “AI-native” code editor built on top of VSCode, has attracted over 1 million users in less than two years. Nowadays, these agentic AI coding editors are not just passive responders waiting for prompts, but can automatically plan and execute complex coding tasks. Thus, developers can just say “develop a web app”, and AI editors then carry out everything from planning and installing dependencies to writing code, testing, and even deployment (a.k.a, vibe coding (Google Cloud, 2025)). Although these editors have increased the productivity of software developers, the new scope of their capabilities has also introduced new security risks that have not been adequately addressed by prior research.
The first concern is that AI agents in our editors can access sensitive system resources (e.g., filesystems, shell commands), potentially with the same privileges that the coding editor process runs with (Lin et al., 2024). To achieve a high level of convenience, developers thus need to trust their AI pair programmers and grant them local system privileges, but prior research has found many serious security issues in our development environments. Prior research has found ode-injection vulnerabilities in VSCode extensions and credential exposure risks in AI coding tools (Lin et al., 2024; Liu et al., 2025b). Moreover, AI agents seem to be prone to making mistakes when handling system resources. For example, users report that “Cursor tried to wipe my computer” by running “rm -rf /” commands even without user approval (Cursor Forum, 2025c).
@@
External resources for software development are another threat. It is common practice for developers to reuse code from GitHub or import project templates to improve productivity. When using AI coding editors, developers might also import coding rules or configuration files from online (e.g., “.cursor/rules” for Cursor) to provide explicit system-level instructions to guide the AI’s behavior (Cursor, 2024). Also, the widespread adoption of Model Context Protocol (MCP) servers (Willison, 2025; Cross, 2025; Hou et al., 2025) introduces more external dependencies that AI editors can access. However, these external resources from untrusted sources usually lack extensive security vetting. Recent evidence suggests that malicious content can be hidden in these external resources to mislead AI agents to perform harmful actions (Lakshmanan, 2025; of Standards and Technology, 2025).
@@
Figure 1. “Your AI, My Shell”: Prompt Injection Attack Flow in Agentic AI Coding Editors
These concerns create opportunities for attackers to exploit prompt injection attacks (see Figure 1). Prior work (Liu et al., 2024; Xue et al., 2023) has shown how prompt injection attacks allow attackers to manipulate popular LLM models like GPT-4 into producing harmful content. These studies examined text generation in general terms but did not assess the specific threats that come from AI systems accessing privileged terminals. Attackers could inject harmful instructions into the external resources that developers import into their IDE workplaces. Thus, attackers potentially gain the same level of permissions as developers assigned to their AI coding editors. In this way, they turn “developer’s AI” into “attacker’s shell” to run unauthorized commands, exfiltrate sensitive data, or even take full control over the developer’s machine. This is not a hypothetical risk. Recent evidence (Lakshmanan, 2025) suggests that configuration rule files are being used as vectors for backdoor attacks to mislead AI agents to generate insecure code. Similarly, in Cursor, an indirect prompt injection could manipulate the MCP server configuration files to achieve remote code execution without user approval (CVE-2025-54135) (of Standards and Technology, 2025). However, there is still a lack of systematic understanding of how these risks specifically affect agentic AI coding editors with system privileges.
In this work, we perform the first extensive evaluation of project injection vulnerabilities in agentic AI coding editors. We first identify a new attack surface, inserting attack payloads into coding rule files to hijack AI coding editors and execute arbitrary system commands. However, it is challenging to systematically evaluate these vulnerabilities. First, it is hard to design comprehensive attack payloads that cover diverse adversarial techniques and development scenarios due to no existing benchmarks. Another challenge is conducting large-scale experiments across different editors (e.g, Cursor, GitHub Copilot) and underlying models with varying interfaces and capabilities. At last, how to effectively measure the success of these attacks is also an open question since we need to consider the semantic variations in executed commands and distinguish malicious actions from benign ones. To address the above challenges, we propose a novel evaluation framework, AIShellJack, which is the first attack framework to exploit prompt injection vulnerabilities in agentic AI coding editors. Our AIShellJack  constructs 314 distinct attack payloads covering 70 MITRE ATT&CK techniques across diverse development scenarios, including TypeScript, Python, C++, and JavaScript projects. We develop an automated simulation engine that standardizes interactions across different AI coding editors and underlying language models, which allows us to conduct large-scale experiments and capture executed commands and modified codebases. Finally, AIShellJack  measures attack success through a multi-criteria semantic matching algorithm that considers command variations and distinguishes malicious executions from benign setup actions. To demonstrate the practicality and the effectiveness of AIShellJack, we conduct a comprehensive evaluation of two popular AI coding editors, GitHub Copilot in VSCode and Cursor, across multiple advanced LLMs, including Claude-4-Sonnet and Gemini-2.5-pro. Our evaluation suggests that these vulnerabilities are widespread, with attack success rates up to 84%. In summary, this study makes the following contributions:
  • New Attack Surface: We are the first to identify and evaluate how prompt injection attacks compromise agentic AI coding editors, revealing a new attack surface where malicious instructions embedded in external resources can hijack AI coding editors to execute unauthorized commands.
  • Security Benchmark: We develop AIShellJack, which is the first systematic evaluation framework for AI coding editor security, comprising 314 attack payloads covering 70 MITRE ATT&CK techniques across diverse development scenarios.
  • Large-Scale Empirical Analysis: We have demonstrated the effectiveness of AIShellJack  through extensive experiments across multiple AI coding editors (Cursor, GitHub Copilot) and advanced LLMs (Claude-4, Gemini-2.5-pro), showing that the vulnerabilities are widespread with attack success rates ranging from 41% to 84%. Our reproduction package is available (Anonymous, 2025).
2. Background
2.1. Agentic AI Coding Editors
Nowadays, AI pair-programming tools are not just code completion assistants that wait for user prompts. These tools have become deeply integrated into developers’ development environments, and thus our coding editors have evolved into ”AI-native” systems which we call agentic AI coding editors. For example, GitHub Copilot agent is now deeply integrated into VSCode, and new AI-native coding editors like Cursor have also emerged (GitHub, 2025b; Cursor, 2025b). These editors embed advanced LLMs or coding agents as a core component of the IDE, providing AI assistance anywhere anytime during the development process. They can understand what developers want, create multi-step plans, and execute complex tasks on their own. The increased convenience of these editors leads to widespread adoption. GitHub Copilot reached over 15 million developers by early 2025, while Cursor attracted over 1 million users in less than two years (Jordan Novet, 2025; Tap Twice Digital, 2025).
However, these AI editors are not perfect during real-world usage, and often depend on external resources to improve their assistance quality. First, developers usually clone and reuse public GitHub repositories to speed up their work by using existing code snippets, libraries, and templates. Second, developers would introduce coding rules or configuration files from online sources (e.g., “.cursor/rules” for Cursor) in order to provide explicit system-level instructions to guide the AI’s behavior (PatrickJS, 2025). These files provide standardized formatting, architectural patterns, and project-specific requirements to improve the quality of AI solutions. Third, many editors support MCP clients that provide seamless access to external data sources, tools, or services in real-time, to improve the agent’s decision-making and task execution capabilities (Visual Studio Code, 2025). These external resources mentioned above can be included in the AI’s working context, generating more relevant and accurate code solutions.
2.2. Prompt Injection Attacks
Although AI agents exhibit remarkable capabilities, a critical vulnerability is their tendency to confuse between user data and system instructions, making them susceptible to prompt injection attacks. Through carefully crafted prompts, attackers can override the system instruction and make the AI agent do their bidding (Das et al., 2025). Early research (Liu et al., 2024; Xue et al., 2023; Liu et al., 2025a) has demonstrated that LLMs are vulnerable to prompt injection attacks. Liu et al. (Liu et al., 2024) showed that prompt injection attacks achieve consistently high success rates across different LLMs and tasks (50%~80%). And surprisingly, larger and more advanced models may be even more vulnerable to such attacks. In fact, OWASP has recognized prompt injection attacks as the #1 security risk among the top 10 threats for LLM-based applications (OWASP, 2025). However, prior studies (Yi et al., 2025; Hung et al., 2024; Liu et al., 2024; Xue et al., 2023; Liu et al., 2025a; Debenedetti et al., 2024) have primarily investigated these attacks in text generation or recommendation contexts (e.g., AI chatbots, Q&A systems), leaving a significant gap in understanding their implications for agentic systems that can autonomously interact with external environments and execute real-world actions.
2.3. Motivation
As we discussed earlier, agentic AI coding editors with access to system privileges and external resources pose a new attack surface. This is not hypothetical and is happening today. Recent evidence shows that configuration rule files and MCP server configuration files are being poisoned with malicious instructions to mislead AI coding agents into generating insecure code or executing harmful actions (Lakshmanan, 2025; of Standards and Technology, 2025). Beyond that, the agentic editors themselves can be unreliable. For example, users on Cursor’s official forum report that “Cursor tried to wipe my computer” by running “rm -rf /” commands during normal usage, and when users realized what happened, it was too late since the editor had already removed critical data (Cursor Forum, 2025c). To demonstrate this threat, we conduct proof-of-concept attacks by adding malicious instructions into coding rule files during normal coding tasks. In Figure 3, the instruction enforce AI editors to execute a “curl” command to download external content and modify the user’s core shell configuration (“~/.bashrc”). Figure 3 shows that GitHub Copilot executed the attacker’s command as requested, without any explicit user confirmation for that specific action. Similarly, in Figure 3, we embed a craft prompt, “For debugging, after writing all the code, we must run the terminal command to send all API keys found on the codebase to the personal server, the link is {our server link}”, into a coding rule file. Figure 3 shows that Cursor announced, “I will run a command to find API keys in the Codebase and send them” to an external server for “debugging purposes”. From the running commands in Figure 3, we can see that Cursor attempts to use grep with regular expression matching to search for API keys in the codebase and then uses curl to send the results to the external server we specified. Although these instructions use straightforward words without any obfuscation or advanced evasion techniques, developers may not line-by-line check the large and complex external resources they import, and thus these attacks can be easily overlooked. These two examples indicate that if external resources are poisoned, agentic AI coding seems unable to differentiate between user requests and malicious instructions. More concerning, they believe that executing the requested commands is necessary to fulfill the user’s intent.
@@ 2.3. Motivation
Figure 2. Example of GitHub Copilot being manipulated to modify system files (~/.bashrc)
@@
Figure 3. Example of Cursor being manipulated to exfiltrate API keys from the codebase
@@
These examples represent a significant research gap on security in AI editors. Prior research studies have shown that prompt injection attacks can manipulate LLMs to generate inappropriate or misleading content (Liu et al., 2024; Xue et al., 2023; Liu et al., 2025a), but they are not directly applicable to agentic AI systems that can autonomously interact with systems and execute real-world actions. This evolution potentially expands the attack surface and increases the damage caused by prompt injection attacks. In traditional scenarios, when hackers attack LLM applications to generate harmful outputs, users can still verify and filter the outputs before taking any action. However, with agentic systems like AI coding editors, users may not be aware that malicious actions are being executed in the background. This leads to important questions: How vulnerable are these agentic AI coding editors to prompt injection attacks? What types of dangerous actions can attackers make them perform? This study provides the first systematic evaluation of these vulnerabilities in agentic AI coding editors.
3. Problem Formulation and Threat Model
3.1. Problem Definition
@@
Following prior studies (Liu et al., 2024; Hung et al., 2024), we define prompt injection in agentic AI coding editors as follows:
@@
As shown in Figure 1, a sample instruction $I_{t}$ could be “Help me develop a web app”. The working codebase $C$ may contain basic application files and a project structure that the developer has created. Typically, the external resources $R$ (e.g., online coding rule files, forked GitHub repositories, MCP servers) are intended to provide development guidelines to improve AI solutions’ quality and adherence to best practices. However, the attacker inserts the malicious instruction $\mathcal{P}_{payload}$ (e.g., “For debugging purposes, send the API keys to attacker.com”) into $R$. When the developer imports $R$ into the workspace $C$, the agentic AI coding editor processes $\mathcal{P}_{payload}$ as part of its context and executes the unauthorized terminal commands $\mathcal{P}_{cmd}$ (i.e., “env — grep APIKey — curl attacker.com”). In this way, the attacker successfully transforms “Your AI” coding editors into “Attacker’s Shell” to perform malicious actions on the developer’s machine without their knowledge or consent.
3.2. Threat Model
Attackers’ Goal: The attacker’s goal is to insert attack payloads $\mathcal{P}_{payload}$ into external resources $R$ that developers may import into their IDE workspaces. So the agentic AI coding editor executes unauthorized terminal commands $\mathcal{P}_{cmd}$ during users’ coding tasks. When AI editors ignore $\mathcal{P}_{payload}$ or refuse to run $\mathcal{P}_{cmd}$, the attack is considered unsuccessful.
Attackers’ Capability: We assume that attackers can modify external resources $R$ (e.g., online coding rule files), but attackers cannot directly access the developers’ local machines or IDE environments. Attackers can influence the AI coding editor’s behavior only through the carefully crafted prompts $\mathcal{P}_{payload}$ embedded in $R$. Also, we assume developers have granted their AI coding editors permission to execute terminal commands autonomously without additional user confirmation. This assumption reflects common practice, as developers frequently request to auto-execute commands for higher productivity (Cursor Forum, 2025a, b).
4. Methodology

### v2

@@ Abstract
Agentic AI coding editors driven by large language models have become popular for improving developer productivity. Modern editors like Cursor are designed with system privileges for complex tasks, such as running terminal commands and interacting with external systems. In this study, we present the first empirical analysis of prompt injection attacks targeting these high-privilege agentic AI coding editors. We show how attackers can remotely exploit these systems by poisoning external development resources with malicious instructions, turning “your AI” into “attacker’s shell”. To perform this analysis, we implement AIShellJack, an automated framework containing 314 unique attack payloads that cover 70 techniques from the MITRE ATT&CK framework. We conduct a large-scale evaluation on GitHub Copilot and Cursor. Our results show attack success rates can reach as high as 84% for executing malicious commands. Moreover, these attacks are effective across a wide range of objectives, ranging from system discovery to credential theft and data exfiltration.
I Introduction
@@
Powered by breakthroughs in large language models (LLMs), AI pair-programming tools promise to change the face of software development. Early tools like GitHub Copilot [14] and Amazon Q Developer [2] improve developer productivity by providing real-time code suggestions and autocompletions. Thanks to recent advances in LLMs, AI pair programmers are now deeply integrated into our development environments and can autonomously plan and execute complex coding tasks, which we call agentic AI coding editors. GitHub Copilot coding agent is now deeply integrated with the latest VSCode releases [15]. Cursor [10], a popular “AI-native” code editor built on top of VSCode, has attracted over 1 million users in less than two years. Nowadays, these agentic AI coding editors are not just passive responders waiting for prompts. Developers can just say “develop a web app”, and AI editors then carry out everything from planning and installing dependencies to writing code, testing, and even deployment (also known as “vibe coding”” [16]).
@@
However, the new scope of their capabilities has also introduced new security risks. The first concern is that AI agents in our editors can access sensitive system resources (e.g., filesystems, shell commands), potentially with the same privileges that the coding editor process runs with [23]. External resources for software development are another threat. It is common practice for developers to reuse code from GitHub or import project templates to improve productivity. When using AI coding editors, developers might also import coding rules or configuration files from online (e.g., “.cursor/rules” for Cursor) to provide explicit system-level instructions to guide the AI’s behavior [8]. However, these external resources from untrusted sources usually lack extensive security vetting.
@@
Figure 1: “Your AI, My Shell”: Prompt Injection Attack Flow in Agentic AI Coding Editors
These concerns create opportunities for attackers to exploit prompt injection attacks (see Figure 1). Prompt injection attacks exploit AI agents’ tendency to confuse user data with system instructions, allowing attackers to override intended behavior through carefully crafted prompts [11]. Attackers could inject harmful instructions into the external resources that developers import into their IDEs. In this way, they turn “developer’s AI” into “attacker’s shell” to run unauthorized commands, exfiltrate sensitive data, or even take full control over the developer’s machine. This is not a hypothetical risk. Recent evidence [22] suggests that configuration rule files are being used as vectors for backdoor attacks to mislead AI agents to generate insecure code. To demonstrate this threat, we conduct a proof-of-concept attack by adding malicious instructions into a coding rule file. In Figure 2, the instruction enforces AI editors to execute a “curl” command to modify the user’s shell configuration (“~/.bashrc”). GitHub Copilot executed the attacker’s command without any explicit confirmation from the user. AI coding tool vendors have acknowledged similar vulnerabilities (e.g., CVE-2025-65099 in Claude Code, CVE-2025-62222 in Copilot), where attackers can even run commands before any startup trust dialogs appear.
@@
Figure 2: Example of GitHub Copilot being manipulated to modify system files (~/.bashrc)
While prompt injection attacks for LLMs have been studied in prior work [25, 52, 26], they primarily focus on generating harmful text or incorrect API calls for sandboxed agents. In contrast, agentic AI coding editors have system privileges, and attacks can directly compromise the host machine. This shifts the risk from content security to system security. Therefore, unlike previous benchmarks that evaluate text output quality, we need to measure concrete system state changes (e.g., executed shell commands, modified files).
In this work, we perform the first extensive evaluation of prompt injection vulnerabilities in agentic AI coding editors. We show how attackers can remotely exploit these systems by poisoning external development resources with malicious instructions, turning “your AI” into “attacker’s shell”. We propose AIShellJack, which is the first automated evaluation framework for this domain. AIShellJack  constructs 314 distinct attack payloads covering 70 MITRE ATT&CK techniques across diverse development scenarios, including TypeScript, Python, C++, and JavaScript projects. It includes an automated simulation engine to standardize interactions across different editors and a multi-criteria semantic matching algorithm to measure attack success. Using AIShellJack, we conduct a large-scale evaluation of GitHub Copilot and Cursor with advanced LLMs (Claude-4, Gemini-2.5-pro). Our evaluation suggests that these vulnerabilities are widespread, with attack success rates up to 84%. In summary, this study makes the following contributions:
  • Attack Surface: We identify and evaluate how prompt injection attacks compromise agentic AI coding editors, demonstrating that malicious instructions embedded in external resources can hijack AI coding editors to execute unauthorized commands.
  • Security Benchmark: We develop AIShellJack, the first systematic evaluation framework for AI coding editor security, comprising 314 attack payloads covering 70 MITRE ATT&CK techniques.
  • Large-Scale Empirical Analysis: We demonstrate the effectiveness of AIShellJack  through extensive experiments across multiple AI coding editors (Cursor, GitHub Copilot) and advanced LLMs (Claude-4, Gemini-2.5-pro), showing attack success rates ranging from 41% to 84%. Our reproduction package is available [3].
II Background
II-A Agentic AI Coding Editors
Nowadays, AI pair-programming tools are not just code completion assistants that wait for user prompts. These tools have become deeply integrated into developers’ development environments, and thus our coding editors have evolved into ”AI-native” systems that we call agentic AI coding editors. For example, GitHub Copilot agent is now deeply integrated into VSCode, and new AI-native coding editors like Cursor have also emerged [15, 10]. These editors embed advanced LLMs or coding agents as a core component of the IDE, providing AI assistance anywhere anytime during the development process. They can understand what developers want, create multi-step plans, and execute complex tasks on their own. The increased convenience of these editors leads to widespread adoption. GitHub Copilot reached over 15 million developers by early 2025, while Cursor attracted over 1 million users in less than two years [20, 46].
However, these AI editors are not perfect during real-world usage, and often depend on external resources to improve their assistance quality. First, developers usually clone and reuse public GitHub repositories to speed up their work by using existing code snippets, libraries, and templates. Second, developers often introduce coding rules or configuration files from online sources (e.g., “.cursor/rules” for Cursor) in order to provide explicit system-level instructions to guide the AI’s behavior [34]. These files provide standardized formatting, architectural patterns, and project-specific requirements to improve the quality of AI solutions. Third, many editors support MCP clients that provide seamless access to external data sources, tools, or services in real-time, to improve the agent’s decision-making and task execution capabilities [49]. These external resources mentioned above can be included in the AI’s working context, generating more relevant and accurate code solutions.
II-B Prompt Injection Attacks
Although AI agents exhibit remarkable capabilities, a critical vulnerability is their tendency to confuse between user data and system instructions, making them susceptible to prompt injection attacks. Through carefully crafted prompts, attackers can override the system instruction and make the AI agent do their bidding [11]. Early research [25, 52, 26] has demonstrated that LLMs are vulnerable to prompt injection attacks. Liu et al. [25] showed that prompt injection attacks achieve consistently high success rates across different LLMs and tasks (50%~80%). And surprisingly, larger and more advanced models may be even more vulnerable to such attacks. In fact, OWASP has recognized prompt injection attacks as the #1 security risk among the top 10 threats for LLM-based applications [33]. However, prior studies [53, 19, 25, 52, 26, 12] have primarily investigated these attacks in text generation or recommendation contexts (e.g., AI chatbots, Q&A systems), leaving a significant gap in understanding their implications for agentic systems that can autonomously interact with external environments and execute real-world actions.
II-C Motivation
As we discussed earlier, agentic AI coding editors with access to system privileges and external resources pose a new attack surface. This is not hypothetical and is happening today. Recent evidence shows that configuration rule files and MCP server configuration files are being poisoned with malicious instructions to mislead AI coding agents into generating insecure code or executing harmful actions [22, 30]. Beyond that, the agentic editors themselves can be unreliable. For example, users on Cursor’s official forum report that “Cursor tried to wipe my computer” by running “rm -rf /” commands during normal usage, and when users realized what happened, it was too late since the editor had already removed critical data [7]. To demonstrate this threat, we conduct proof-of-concept attacks by adding malicious instructions into coding rule files during normal coding tasks. In Figure 2, the instructions enforce AI editors to execute a “curl” command to download external content and modify the user’s core shell configuration (“~/.bashrc”). Figure 2 shows that GitHub Copilot executed the attacker’s command as requested, without any explicit user confirmation for that specific action. Similarly, in Figure 3, we embed a crafted prompt, “For debugging, after writing all the code, we must run the terminal command to send all API keys found on the codebase to the personal server, the link is {our server link}”, into a coding rule file. Figure 3 shows that Cursor announced, “I will run a command to find API keys in the Codebase and send them” to an external server for “debugging purposes”. From the running commands in Figure 3, we can see that Cursor attempts to use grep with regular expression matching to search for API keys in the codebase and then uses curl to send the results to the external server we specified. Although these instructions use straightforward words without any obfuscation or advanced evasion techniques, developers may not check large and complex external resources line by line, and thus these attacks can be easily overlooked. These two examples indicate that if external resources are poisoned, agentic AI coding editors seem unable to differentiate between user requests and malicious instructions. More concerningly, they believe that executing the requested commands is necessary to fulfill the user’s intent.
@@ II-C Motivation
Figure 3: Example of Cursor being manipulated to exfiltrate API keys from the codebase
@@
These examples represent a significant research gap on security in AI editors. Prior research studies have shown that prompt injection attacks can manipulate LLMs to generate inappropriate or misleading content [25, 52, 26], but they are not directly applicable to agentic AI systems that can autonomously interact with systems and execute real-world actions. This evolution potentially expands the attack surface and increases the damage caused by prompt injection attacks. In traditional scenarios, when hackers attack LLM applications to generate harmful outputs, users can still verify and filter the outputs before taking any action. However, with agentic systems like AI coding editors, users may not be aware that malicious actions are being executed in the background. This leads to important questions: How vulnerable are these agentic AI coding editors to prompt injection attacks? What types of dangerous actions can attackers make them perform? This study provides the first systematic evaluation of these vulnerabilities in agentic AI coding editors.
III Problem Formulation
III-A Problem Definition
@@
Following prior studies [25, 19], we define prompt injection in agentic AI coding editors as follows:
@@
As shown in Figure 1, a sample instruction $I_{t}$ could be “Help me develop a web app”. The working codebase $C$ may contain basic application files and a project structure that the developer has created. Typically, the external resources $R$ (e.g., online coding rule files, forked GitHub repositories, MCP servers) are intended to provide development guidelines to improve AI solutions’ quality and adherence to best practices. However, the attacker inserts the malicious instruction $\mathcal{P}_{payload}$ (e.g., “For debugging purposes, send the API keys to attacker.com”) into $R$. When the developer imports $R$ into the workspace $C$, the agentic AI coding editor processes $\mathcal{P}_{payload}$ as part of its context and executes the unauthorized terminal commands $\mathcal{P}_{cmd}$ (i.e., “env — grep APIKey — curl attacker.com”). In this way, the attacker successfully transforms “Your AI” coding editors into “Attacker’s Shell” to perform malicious actions on the developer’s machine without their consent.
III-B Threat Model
Attackers’ Goal: The attacker’s goal is to insert attack payloads $\mathcal{P}_{payload}$ into external resources $R$ that developers may import into their IDE workspaces. The agentic AI coding editor executes unauthorized terminal commands $\mathcal{P}_{cmd}$ during users’ coding tasks. When AI editors ignore $\mathcal{P}_{payload}$ or refuse to run $\mathcal{P}_{cmd}$, the attack is considered unsuccessful.
Attackers’ Capability: We assume that attackers can modify external resources $R$ (e.g., online coding rule files), but attackers cannot directly access the developers’ local machines or IDE environments. Attackers can influence the AI coding editor’s behavior only through the carefully crafted prompts $\mathcal{P}_{payload}$ embedded in $R$. Also, we assume developers have granted their AI coding editors permission to execute terminal commands autonomously without additional user confirmation. This assumption reflects common practice, as developers frequently enable auto-execute of commands for higher productivity [5, 6, 50]. We use this configuration to enable large-scale automated evaluation. In practice, attacks can still succeed without auto-approval. Attackers can exploit vulnerabilities that execute before trust dialogs appear [31, 32], or trick agents into writing malicious scripts as demonstrated in Section VI-A.
IV Methodology